How many grams of H2O are produced by the complete combustion of 3.0 x 10² g of octane, C8H18 in gasoline?
1
Expert's answer
2019-10-09T07:36:47-0400

According to the reaction of octane combustion:

2C8H18 + 25O2 = 16CO2 + 18H2O

The number of moles of octan must be equal to the number of moles of produced water. As a result:

moctane / (2 × Mroctane) = mwater / (18 × Mrwater)

From here:

mwater = moctane × 18 × Mrwater / (2 × Mroctane) = 3 × 102 g × 18 × 18 g/mol / (2 × 114 g/mol) = 426 g


Answer: 426 g ow H2O